OpenStudy (loser66):

If \(X^2 =I_3\) what is X? matrix problem Please, help

OpenStudy (loser66):

I know how to find X by diagonal matrix except that method, is there any other way to find out X ?

OpenStudy (anonymous):

the other alternative is through intuition... if \(X^2=I\) then we know it could be a reflection or a rotation by \((2n+1)\pi\) (since then \(2(2n+1)\pi=(2n+1)2\pi\) is therefore a nil rotation)

OpenStudy (anonymous):

now...what combinations of these could we use to make more solutions

OpenStudy (loser66):

Is it.... Chinese!! hihihi, but I can learn, Please teach me,